A uniform charge wire, linear charge density d, is laid in the form of a semi circle of radius r. Find the electric field generated by the semi circle of the center


1
Expert's answer
2021-10-07T09:21:08-0400

The electric field is


"E_x=\frac{kd}{r}\int^\pi_0\cos\theta\text d\theta=\frac{kd}r\Big[\sin\theta\Big]^\pi_0=0,\\\space\\
E_y=\frac{kd}{r}\int^\pi_0\sin\theta\text d\theta=\frac{kd}r\Big[\cos\theta\Big]^\pi_0=-\frac{2kd}r."

Ex is the field along the diameter, Ey is the field along the radius of the semi-circle (vertical radius perpendicular to the base of the semi-circle if we put it branches down).